OpenStudy (anonymous):

The solutions to the inequalities 0≤x≤2, 0≤y≤1, and y≤−1/ 2 x+3/ 2 form a polygonal region in the plane.? how many sides does this polygon have?

OpenStudy (anonymous):

i need help please!

OpenStudy (anonymous):

0≤x≤2, 0≤y≤1, forms a rectangle with corners : O(0,0) , A(2,0) , B(2,1) , J(0,1) and y = -(1/ 2) x + (3/ 2) passes through : C(0, 3/2) and D(3,0) ---> line (L) and (L) cuts the rectangle at two points : E(1,1) so that y = 1 = -(1/2)x+ 3/2 gives: x = 1 and F(2, f ) so that : f = -(1/ 2) 2 + (3/ 2) = 1/2 therefore : the polygon has 5 sides

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

helpful but i dont get the part you get 4 sides?

OpenStudy (anonymous):

where from?

OpenStudy (anonymous):

\(0\le x\le2\):